Calculate the pH of a solution containing 0.70 M HNO2 and 0.40 M NaNO2. The K, for HNO2 is 4.0x10-4
Answer #1

using H-H equation, pH = pka + log[Na Nor] [HNO2) = -log ka + log (NaN a NO2 --dog (9.0x164) + 1096o.) = -43.39794) + 60.2430